OpenStudy (anonymous):

y = e^(2x)^13 what is the derivative, posted it on wolfram, still couldnt get it right

OpenStudy (accessdenied):

The function is y = e^((2x)^13) ?

OpenStudy (accessdenied):

I'm thinking we would do some chain rule here, with f(x) = e^x being the outer function and g(x) = (2x)^13 being the inner function.

OpenStudy (anonymous):

Yeah that is the problem, and yeah I think it is chain rule

OpenStudy (bahrom7893):

e^(2^13*x^13) right?

OpenStudy (bahrom7893):

(e^u)' = e^u * u' [ e^(2^13*x^13) ]' = e^(2^13*x^13) * 2^13 * 13x^12

OpenStudy (anonymous):

Idk, This is so confusing to me, I think I have two tries left, what do yall think?

OpenStudy (bahrom7893):

webworks?

OpenStudy (bahrom7893):

can u post a screenshot?

OpenStudy (anonymous):

\[e ^{2x}^{13}\]

OpenStudy (anonymous):

what does that look like?

OpenStudy (bahrom7893):

woah wait.. this is soo simple.. im a dumbass

OpenStudy (bahrom7893):

(a^b)^c = a^(bc) e ^{2x}^{13} = e^{26x} => derivative: 26*e^(26x)

OpenStudy (anonymous):

my homework is so stupid it says its wrong, im asking my teacher tomorrow!